# LLM-Based Intent Processing and Network Optimization Using Attention-Based Hierarchical Reinforcement Learning

### v1

@@ IV-B Intent Validation Using Transformer
To perform intent validation, we utilize one of our previous works [4] where traffic data are collected so that it can be fed to the transformer-based time series predictor for future traffic volume prediction. The traffic predictor resides in the non-RT-RIC as an rApp (named as intent validation rApp, as in Fig 2.). The forecast for traffic volume is performed for successive time intervals, with two distinct thresholds derived from historical data. The higher threshold, denoted as $Th_{p}$, is set based on previous instances of heavy traffic. If the forecast traffic surpasses $Th_{p}$, RIC applications such as traffic steering and power allocation can be activated to enhance or sustain high throughput, depending on the intent of an MNO. The lower threshold, $Th_{t}$, is established based on low traffic volume. In the event that the anticipated traffic decreases to less than this threshold, an intent on increasing energy efficiency can be validated leading to the initiation of energy-saving rApp like cell sleeping due to the expected lower traffic levels. The algorithmic process of predictive traffic validation is presented in Algorithm 1 for the overall representation of the process. Note that the intent validation can also be performed by predicting other crucial network parameters or values like congestion levels, device connectivity, status, and so on.
@@
1:Predict future traffic $T_{p}$ for the upcoming time $t+1$ using Autoformer
2:Extract intent ($I_{MNO}$) details: Action (increase/decrease), Metric $c$, and Change Percentage (%) using LLM
3:Define traffic classes $K=\{k_{1},k_{2},\ldots,k_{i}\}$
4:QoS profile of a traffic class: $Q_{s}=\{Q_{s_{1}},Q_{s_{2}},\ldots,Q_{s_{j}}\}$
5:if ($T_{p}>Th_{p}$ ||||| | $T_{p}<Th_{t}$then
6:     Check QoS profiles for the relevant traffic class $k$
7:     if $I_{MNO}$ impacts $Q_{s}$ then
8:         for each QoS parameter $Q_{s_{j}}$ in $k_{i}$’s profile do
9:              Calculate QoS drift $I_{Q}(Q_{s_{j}})$ using eq. 2
10:              if $I_{Q}(Q_{s_{j}})\neq 0$ then
11:                  Intent invalid
12:                  break
13:              end if
14:         end for
15:         if $I_{Q}$ for elements in $Q_{s}$ is 0 ||||| | $A_{QoS}>D_{QoS}$ then
16:              Validate intent
17:         end if
@@
19:else
20:     Re-calculate $Th_{p}$ and $Th_{t}$
21:end if
22:return Validation results
@@ IV-C Network Optimization Using Attention-based HRL
Once the intent is validated, it allows us to go to the final stage of this work which is optimizing the network performance based on the operator intent using an RL algorithm. To do that, we opt for hierarchical Deep-Q-Network (h-DQN), an HRL algorithm having a two-level hierarchy with a meta-controller on top [12]. Meta-controller can take in a network state (e.g., traffic class) and a goal (desired change of a performance metric extracted from an intent) to achieve and the controller in the lower level takes the action of choosing an RIC application or combinations of them based on the state and the goal. The meta-controller is designed as an rApp in the non-RT-RIC (see Fig. 2). RIC applications are controlled by the meta-controller that can directly optimize system performance are summarized as follows:
IV-C1 Traffic Steering xApp
The traffic steering xApp is engineered to concurrently uphold QoS across various traffic types by employing a steering mechanism utilizing a DQN. To ensure optimal performance, the xApp’s state and reward functions are designed with a focus on two KPIs: network latency and overall system throughput. For more detailed insights into this xApp, please refer to [13].
IV-C2 Cell Sleeping rApp
The cell sleeping rApp is based on DQN and aims to reduce network power consumption by deactivating idle or less busy BSs based on traffic load and queue length (set of states). Higher rewards are provided for taking actions that increase energy efficiency without overloading the active BSs. Technical details of this DQN-based rApp can be found in our previous work [7].
IV-C3 Beamforming xApp
Similar to the past two apps, we also employ DQN to develop the beamforming xApp where UE coordinates are used as a set of states. The set of actions comprises $\alpha(\chi_{n})$ and $\delta_{n}$. Here, $\chi$ represents the steering angle, and $\alpha(\chi_{n})$ denotes the array steering vector corresponding to the direction $\chi_{n}$ of the $n^{th}$ element in the codebook. $\delta_{n}$ signifies the adjustment in power level. The selection of optimal actions is determined by a reward function composed of the sum of two components: i) the ratio of the system’s achieved throughput to the specified throughput requirement for a given traffic type $k$, and ii) the ratio of the energy efficiency linked with the base station’s throughput and transmission power to the maximum theoretical energy efficiency.
IV-C4 Power Allocation xApp
The power allocation xApp tries to maximize the total throughput. The set of states includes transmission rate and channel state information. We divide the transmission power levels into $P_{L}$. The action is to choose a power level for each RBG of each BS. $A=\{P_{min},P_{L_{1}},...,P_{L_{2}},P_{max}\}$. The goal is to improve throughput. The reward function is formulated to increase the transmission rate of a BS on a Resource Block Group (RBG).
IV-C5 Handover decision making xApp for energy saving
This xApp is specifically tailored with DQN to achieve energy efficiency via optimally tailored handover policies. To accommodate the extra processing demands and signaling exchanges involved in handover executions at both serving and target base stations, we establish a handover energy consumption function to incorporate the additional energy overheads:
$E_{u}^{HO}(t)=\begin{cases}E^{HO}(t)&\text{if handover occurs at time $t$}\\ 0&\text{otherwise,}\\ \end{cases}$ (3)
where $E^{HO}\geq 0$ denotes the energy consumption at both the serving and target BS during handover execution. The total energy consumption of a serving BS, $b$ at time slot $t$ for a UE, $u$ is expressed as follows: $E_{total}=E_{u,b}^{Tx}+E^{HO}$, where $E_{u,b}$ is the transmission energy consumption when a $u$ is served by a $b$. The objective of this xApp is to reduce average energy consumption. More details on this xApp can be found in [14].
At this stage of the work, we define an MDP to solve the problem formulated in eq.1.
@@ IV-C5 Handover decision making xApp for energy saving
  • State and Action: The set of states consists of traffic flow types of different users in the network. $S=\{T_{1},..,T_{2},...,T_{3},...,T_{4},..\}$. Both meta-controller and controller share the same states. Selecting RIC application, or combinations of them is considered as actions to be performed by the controller which is defined as {$A=A_{App1},A_{App1,2},....,A_{AppN}\}$.
@@
  • Intrinsic reward: The intrinsic reward function ($r_{in}$) for the controller is: $r_{in}=C_{\rho}-\varpi\gamma_{q_{s}}$.
@@
The process of selecting RIC applications for network optimization using HRL can be summarized as follows:
  • Step 1: An input is provided by the MNO associated with a performance metric.
  • Step 2: Intent is processed and crucial information like Action: “Increase/Decrease”. “Which metric” and by how much (in percentage) are extracted using pre-trained ALBERT.
  • Step 3: Extracted information is passed to the intent validation rApp.
  • Step 4: With a valid intent, a target associated with the performance metrics (goals) and feasible action set for the intent are provided to the controller in near-RT-RIC by the meta controller rApp in the non-RT-RIC.
  • Step 5: The controller selects an RIC application or a combination of them to reach the target performance as close as possible.
  • Step 6: Selected RIC applications optimize the performance of the network as a response to the intent of the operator.
We employ an attention mechanism [5] to reduce the action space for our h-DQN agent, enhancing efficiency and decision-making while expediting convergence. A critical component of this process is the Markovian option, $\omega$, which includes a policy $\pi$, a termination condition $\beta$, and an initiation set $I_{s}$. Actions are chosen based on $\pi$ until $\beta$ terminates the option. Our approach uses hard attention [5], selecting a specific part of the input for focus. When an option $\omega$ pursues a goal $g$, the intent completion function assesses if the option’s policy $\pi_{\omega}$ achieves a desired goal within a trajectory, returning a binary outcome. This outcome helps us prioritize affordance-aware options. Multiple options may meet an intent up to a threshold $\epsilon$, and using values of $\epsilon$ and the indicator function $I_{\omega}$, we can efficiently narrow down the action space with a supervised learning algorithm.
@@ V-A Simulation setup
Figure 5: Performance analysis of the proposed method: (a) energy efficiency, (b) network delay, and (c) throughput.
The simulation environment in this work consists of one macro-cell surrounded by a couple of densely deployed small cells having a multi-RAT environment with 60 users in total. For 5G NR, the bandwidth is set at $60$ MHz, with carrier frequencies of $3.5$ GHz and $30$ GHz, subcarrier spacing of $15$ KHz, and maximum transmission power of $43$ dBm, following the 3GPP urban micro channel model [15]. LTE settings include a $40$ MHz bandwidth, $800$ MHz carrier frequency, $15$ KHz subcarrier spacing, and a maximum transmission power of $38$ dBm, modeled after the 3GPP urban macro channel environment [15]. There are four classes of traffic: video, gaming, voice, and Ultra Reliable Low Latency Communications (URLLC) traffic. The traffic characteristics encompass inter-arrival times between packets, $12.5$, $40$, $20$, and $0.5$ ms respectively, distributed according to Pareto (video), Uniform (gaming), and Poisson distributions (voice and URLLC). The packet sizes are set to be $250$, $120$, $60$, and $32$ bytes, corresponding to the specified traffic classes. These values are fixed based on [16]. We refer to traffic steering, cell sleeping, beamforming, handover management, and power allocation applications as xApp1, rApp2, xApp3, xApp4, and xApp5 respectively in the next section to present our results.
@@ V-B Simulation Results
The first simulation result is associated with intent validation. When there is no intent validation, conflicting intents from an operator can severely degrade performance. Fig. 3 presents such an example where we have a low traffic load of 5Mbps. At $2200^{th}$ time slot, there is a sharp decrease in energy efficiency. This is because of the unwanted intent of increasing throughput even though the cell sleeping application is active due to low traffic demand. An intent of increasing throughput by an MNO would result in activating more cells even though not necessary and decrease overall energy efficiency. Similar impacts can be observed for other KPIs like delay, and throughput also.
As shown in Fig. 4, the operator’s goal to “increase throughput” leads to specific RIC application selections. For instance, aiming for a $5\%$ throughput increase results in a noticeable surge after several time slots due to xApp1. Aiming for a $10\%$ increase involves both xApp1 and xApp3. Conversely, reducing power consumption significantly decreases throughput around the $1250^{th}$ time slot, reflecting the deactivation of xApp1 and xApp3 and the activation of rApp2. Aiming for a $20\%$ increase, xApp1, xApp3, and xApp4 together produce a sharp throughput rise.
The proposed method’s effectiveness is compared with two baselines: an HRL-based RIC application initiation without an attention mechanism or intent validation, and a single DRL-based application scenario. The proposed method employs a hard attention mechanism, focusing on the subset of RIC applications most relevant for the current traffic type and operator’s intent. This focus reduces conflicting actions and shows superior performance over the HRL baseline (improvements of $12.02\%$ in throughput, $26.5\%$ in delay, and $17.1\%$ in energy efficiency) and the DRL baseline (improvements of $17.25\%$, $48.6\%$, and $39.3\%$ in the same metrics), as observed in Fig. 5.
@@ VI Conclusions
In this paper, first, we process intents using an LLM and extract crucial information regarding which performance metric to optimize and by how much. Next, a Transformer-based intent validation technique has been used to rule out intents that conflict with the contemporary network state to avoid performance degradation. Lastly, we use an attention-based HRL to initiate and orchestrate RIC applications for performance optimization. The proposed method outperforms the HRL baseline in throughput, delay, and energy efficiency by $12.02\%$, $26.5\%$, and $17.1\%$ respectively. Additionally, compared to the DRL baseline, it achieves superior results across the same performance metrics, with enhancements of $17.25\%$, $48.6\%$, and $39.3\%$ respectively. In the future, we plan to perform highly complex intent translation from the MNO and perform intent validation predicting other parameters.
@@ Acknowledgement
This work has been supported by MITACS, Ericsson Canada, and NSERC Canada Research Chairs program.

### v2

@@ IV-B Predictive Intent Validation Using Transformer
To perform intent validation, we utilize one of our previous works [5] where traffic data are collected so that it can be fed to the transformer-based time series predictor for future traffic volume prediction. The traffic predictor resides in the strategic controller (named as intent validation app, as in Fig 1.). The forecast for traffic volume is performed for successive time intervals, with two distinct thresholds derived from historical data. The higher threshold, denoted as $Th_{p}$, is decided based on previous instances of heavy traffic. If the forecast traffic surpasses $Th_{p}$, applications like traffic steering and power allocation can be activated to enhance or sustain high throughput, depending on the intent of a human operator. The lower threshold, $Th_{t}$, is established based on low traffic volume. In the event that the anticipated traffic decreases to less than this threshold, an intent to increase energy efficiency can be validated. It will then lead to the initiation of energy-saving applications like cell sleeping due to the expected lower traffic levels. The algorithmic process of predictive traffic validation is presented in Algorithm 2.
@@
1:Predict traffic $T_{p}$ for $t+1$ and extract $i$ using Algorithm 1
2:if ($T_{p}>Th_{p}$ ||||| | $T_{p}<Th_{t}$then
3:     Check QoS profiles for the relevant traffic class $k$
4:     if $i$ impacts $Q_{s}$ then
5:         for each QoS parameter $Q_{s_{j}}$ in $k_{\eta}$’s profile do
6:              Calculate QoS drift $I_{Q}(Q_{s_{j}})$ using eq. 2
7:              if $I_{Q}(Q_{s_{j}})\neq 0$ then
8:                  Intent invalid
9:                  break
10:              end if
11:         end for
12:         if $I_{Q}$ for elements in $Q_{s}$ is 0 ||||| | $A_{QoS}>D_{QoS}$ then
13:              Validate intent
14:         end if
15:     end if
16:else
17:     Re-calculate $Th_{p}$ and $Th_{t}$
@@
19:return Validation results
@@ IV-C Network Optimization Using Attention-based HRL
After intent validation, we optimize the network performance based on the operator intent using an HRL algorithm. We opt for a hierarchical Deep-Q-Network (h-DQN) having a two-level hierarchy with a meta-controller on top [12] and a controller on the bottom. These two controllers are solely related to the h-DQN algorithm that we are proposing to use to manage network-optimizing applications. Meta-controller and the controller are hosted in the strategic and tactical controller, respectively. We employ an attention mechanism [6] to reduce the action space for our h-DQN agent, enhancing efficiency and decision-making while expediting convergence. We can define the attention function as follows:
$\alpha(s_{t},i,a)=f_{\theta}(s_{t},i,a),$ (3)
where $f_{\theta}$ computes the attention score trained using supervised learning, to predict which actions are most likely to fulfill the operator’s intent. Once the attention scores $\alpha(s_{t},i,a)$ are computed, the action space is reduced to a subset, $A_{s}^{att}$ containing only the most relevant actions.
$A_{s}^{att}=\{a\in A\mid\alpha(s_{t},i,a)>\epsilon\},$ (4)
where $\epsilon$ is a threshold that filters out actions with low relevance. This subset $A_{s}^{att}$ is much smaller than the full action space $A$, allowing the agent to make decisions more efficiently. For each Markovian option $\omega$ [6], the attention mechanism selects only the relevant options based on the operator’s intent. Each option is defined by a policy $\pi_{\omega}$, a termination condition $\beta_{\omega}$, and an initiation set $I_{s}$. After reducing the action space, the agent applies its policy $\pi$ to select the best action from $A_{s}^{att}$:
$\pi(s_{t})=\arg\max_{a\in A_{s}^{att}}Q(s_{t},a),$ (5)
where $Q(s_{t},a)$ is the action-value function that represents the expected return from taking action $a$ in state $s_{t}$.
@@
Meta-controller in h-DQN can take in a network state (e.g., traffic class) and a goal (desired change of a performance metric extracted from an intent) to achieve. The controller in the lower level takes the action of choosing an application or combinations of them based on the state and the goal. The meta-controller is designed as a top-level application in the strategic controller (see Fig. 1). Applications are controlled by the meta-controller that can directly optimize system performance. At this stage of the work, we define the following MDP for the h-DQN to solve the problem formulated in (1).
@@
  • State and Action: The set of states consists of traffic flow types of different users in the network. $S=\{T_{1},..,T_{2},...,T_{3},...,T_{4},..\}$. Both meta-controller and controller share the same states. Selecting network optimizing applications, or combinations of them is considered as actions to be performed by the controller which is defined as {$A=A_{App1},A_{App1,2},....,A_{AppN}\}$.
@@
  • Intrinsic reward: The intrinsic reward function ($r_{in}$) for the controller is: $r_{in}=C_{\rho}-\varpi\gamma_{q_{s}}$ (see (1) for more details.).
@@
We have used five applications in total that are controlled by the meta-controller. They are traffic steering [13], cell sleeping [3], beamforming [3], power allocation [14], and energy efficient handover management [15]. All of these applications are based on DRL. The traffic steering application is engineered to concurrently uphold QoS across various traffic types by employing a steering mechanism to ensure optimal performance in terms of delay and throughput. The cell sleeping app (works on a longer time scale in the strategic controller) aims to reduce network power consumption by deactivating idle or less busy BSs based on traffic load and queue length. The beamforming application uses UE coordinates to sort out the steering angle, and the array steering vector corresponding to the respective element in the codebook. The power allocation application tries to maximize the total throughput and lastly, the handover management app is specifically tailored with DRL to achieve energy efficiency via optimally tailored handover policies. The process of selecting and orchestrating apps for network optimization using HRL can be summarized as follows:
  • Step 1: An input is provided by the human network operator associated with a performance metric.
  • Step 2: Intent is processed and crucial information like intent type: “energy-efficiency/ throughput/ Delay”, and magnitude of change (in percentage) are extracted using pre-trained LLM.
  • Step 3: Extracted information is passed to the intent validation application.
  • Step 4: With a valid intent, a target associated with the performance metrics (goals) and feasible action set for the intent are provided to the controller by the meta-controller.
  • Step 5: The controller selects an application or a combination of them to reach the target performance as close as possible.
  • Step 6: Selected applications optimize the performance of the network as a response to the intent of the operator.
@@ V-A Simulation setup
The simulation setup for this study features a macro cell surrounded by densely deployed small cells in a multi-RAT (multi-radio access technology) environment, accommodating a total of 60 users. For the 5G NR configuration, a bandwidth of 60 MHz is used, with carrier frequencies of 3.5 GHz and 30 GHz, a subcarrier spacing of 15 KHz, and a maximum transmission power of 43 dBm, based on the 3GPP urban micro channel model [16]. The LTE setup, on the other hand, employs a 40 MHz bandwidth, an 800 MHz carrier frequency, a 15 KHz subcarrier spacing, and a maximum transmission power of 38 dBm, following the 3GPP urban macro channel model [16].
There are four traffic types: video, gaming, voice, and Ultra Reliable Low Latency Communications (URLLC), each characterized by specific inter-arrival times for packets: 12.5 ms, 40 ms, 20 ms, and 0.5 ms, respectively. These packet inter-arrival times follow different distributions—Pareto for video, Uniform for gaming, and Poisson for both voice and URLLC. In the following section, we present results using shorthand identifiers for applications: App1 (traffic steering), App2 (cell sleeping), App3 (beamforming), App4 (power allocation), and App5 (handover management).
Table I provides a summary of the tunable parameters for both the Autoformer (used for traffic prediction) and the RL methods (including the HRL control algorithm and DRL-based network applications).
TABLE I: Parameter settings for Autoformer and RL
Autoformer
Learning rate $(\alpha)$, Batch size $10^{-4}$, $32$ [5]
Model architecture 2 encoder layers, 1 decoder layer
HRL parameters (h-DQN)
Exploration Probability ($\epsilon_{2}$) $1$, annealed to $0.1$ [12]
Learning rate ($\alpha$), discount factor $\gamma$ $0.00025$, $0.99$ [12]
RL paramters (applications)
Batch size, Initial exploring steps $32$, $3000$
Learning rate ($\alpha$), discount factor $\gamma$ $0.5$, $0.9$
@@ V-B Simulation Results
The first simulation result is associated with intent validation. When there is no intent validation, conflicting intents from an operator can severely degrade performance. Fig. 2 presents such an example where we have a traffic load of 5Mbps. At $2200^{th}$ time slot, there is a sharp decrease in energy efficiency. This is because of the unwanted intent of increasing throughput even though the cell sleeping application is active due to low traffic demand. An intent of increasing throughput would result in activating more cells unnecessarily and decrease overall energy efficiency. Similar impacts can be observed for other KPIs like delay, and throughput also.
As shown in Fig. 3, the human operator’s goal to “increase throughput” leads to multiple application selection. For instance, aiming for a $5\%$ throughput increase results in a noticeable surge after several time slots due to App1. Aiming for a $10\%$ increase involves both App1 and App3. Conversely, reducing power consumption significantly decreases throughput around the $1250^{th}$ time slot, reflecting the deactivation of App1 and App3 and the activation of App2. When the operator aims for a $20\%$ increase, App1, App3, and App4 together produce a sharp throughput rise.
The proposed method’s effectiveness is compared with two baselines: an HRL-based application initiation without an attention mechanism or intent validation, and a single DRL-based application scenario. The proposed method employs a hard attention mechanism, focusing on the subset of applications most relevant for the current traffic type and operator’s intent. This focus reduces conflicting actions and shows superior performance over the HRL baseline (improvements of $12.02\%$ in throughput, $26.5\%$ in delay, and $17.1\%$ in energy efficiency) and the DRL baseline (improvements of $17.25\%$, $48.6\%$, and $39.3\%$ in the same metrics), as observed in Fig. 4.
Figure 4: Performance analysis of the proposed method: (a) network delay, (b) throughput, and (c) energy efficiency.
@@ VI Conclusions
In this paper, first, we process intents using an LLM and extract crucial information regarding which performance metric to optimize and by how much. Next, a Transformer-based intent validation technique has been used to rule out intents that conflict with the contemporary network state to avoid performance degradation. Lastly, we use an attention-based HRL to initiate and orchestrate applications for performance optimization. The proposed method outperforms the HRL baseline in throughput, delay, and energy efficiency by $12.02\%$, $26.5\%$, and $17.1\%$ respectively. Additionally, compared to the DRL baseline, it achieves superior results across the same performance metrics, with enhancements of $17.25\%$, $48.6\%$, and $39.3\%$ respectively. In the future, we plan to perform highly complex intent translations from the human network operator.
@@ Acknowledgement
This work has been supported by MITACS, Ericsson Canada, and the NSERC Canada Research Chairs program.